A compound contains 0.24 g C and 0.050 g H. The molar mass of the compound is 58 g/mole. What are the empirical and molecular formulas of the compound?
Answer #1

      Element                  mass                A.wt             no of moles

C 0.24g                 12                0.24/12   = 0.02moles

H    0.05g                 1                 0.05/1     = 0.05 moles

      C : H     = 0.02 : 0.05

                    = 0.02/0.02 : 0.05/0.02

                    = 1:2.5

                    = 2: 5

empirical formula = C2H5

molecular formula = (empirical formula)n

                     n          = M.wt/E.F.Wt    = 58/29   = 2

molecular formula = (C2H5)2 = C4H10